OpenStudy (anonymous):

What is sin(13pi/4)? Explain your thinking. I know the answer is sqrt 2/2 , i just want to make sure that it's right.

jimthompson5910 (jim_thompson5910):

it's incorrect, but you're close

jimthompson5910 (jim_thompson5910):

13pi/4 - 2pi = 13pi/4 - 8pi/4 = 5pi/4 sin(13pi/4) = sin(5pi/4) = ??

OpenStudy (anonymous):

i got 4pi/4

OpenStudy (anonymous):

unless i'm doing it wrong @jim_thompson5910

jimthompson5910 (jim_thompson5910):

use this unit circle http://etc.usf.edu/clipart/43200/43215/unit-circle7_43215_lg.gif

OpenStudy (anonymous):

oh! its just negative @jim_thompson5910

OpenStudy (anonymous):

_ sqrt 2/2.

jimthompson5910 (jim_thompson5910):

yes, \[\Large \sin\left(\frac{5\pi}{4}\right) = -\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

yay!! Thank you so much!:) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

You can always do the calculation keeping in mind that \[\pi = 180 \ degrees\]